# Regret Analysis for Randomized Gaussian Process Upper Confidence Bound

### v1

@@ 3 Algorithm
Algorithm 1 shows a pseudo-code of IRGP-UCB. The difference from the original GP-UCB is that $\zeta_{t}$ is a random variable, not a constant. Using generated $\zeta_{t}$, the next input is chosen as follows:
@@
where $\lambda$ is a rate parameter. This can be seen as a distribution of the sum of $s_{t}$ and $Z\sim{\rm Exp}(\lambda)$. Thus, the sampling from the shifted exponential distribution can be easily performed using an exponential distribution as follows:
@@ 4 Regret Analysis
In this section, we analyze the performance of IRGP-UCB through regret analyses. First, we analyze the average-case performance through expected regret, where the expectation is taken with all the randomness, that is, $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and $\{\zeta_{t}\}_{t\geq 1}$. Regarding the expected regret, IRGP-UCB achieves sub-linear upper bounds without increasing the confidence parameter when the input domain is finite. Finally, we analyze the conditional expected regret, where the expectation is taken with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$ and the probability is taken with $\{\zeta_{t}\}_{t\geq 1}$. This analysis shows that IRGP-UCB achieves the sub-linear upper bounds without increasing the confidence parameter when the input domain is finite, even if the expectation is taken only with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$. Lastly, we discuss the high probability regret bounds of the randomized GP-UCB, where the probability is evaluated regarding all the randomness.
@@ 4.1 Expected Regret Bounds
Here, we show the expected cumulative regret bound, which is also called Bayesian cumulative regret (BCR) (Russo and Van Roy, 2014). The definition of BCR is
@@
Suppose that $f$ is a sample path from a GP with zero mean and a positive semidefinite kernel $k$ and ${\cal X}$ is finite. Pick $\delta\in(0,1)$ and $t\geq 1$. Then, for any given ${\cal D}_{t-1}$,
@@
See Appendix A.1 for the proof. This lemma differs slightly from (Srinivas et al., 2010, Lemma 5.1), since, in Lemma 4.1, the iteration $t$ is fixed, and $\beta_{\delta}$ does not depend on $t$.
Based on Lemma 4.1, we show the following lemma, which improves the BCR bound:
@@
We here show the short proof of Lemma 4.2 although detailed proof is shown in Appendix A.2. From Lemma 4.1, we can obtain the following inequality:
@@
Lemma 4.2 shows a direct upper bound of the expectation of $f(\boldsymbol{x}^{*})$. A remarkable consequence of Lemma 4.2 is an upper bound of BCR as follows:
@@
where $v_{t}(\boldsymbol{x})\coloneqq\mu_{t-1}(\boldsymbol{x})+\zeta^{1/2}_{t}\sigma_{t-1}(\boldsymbol{x})$. Importantly, $\mathbb{E}[\zeta_{t}]=2+s_{t}$ is a constant since $s_{t}$ and $\lambda$ for $\zeta$ does not depend on $t$ in Lemma 4.2. Thus, using Lemma 4.2, we can obtain the BCR bound for the finite input domain:
@@
Theorem 4.1 has two important implications. First, Theorem 4.1 shows the convergence rate $O(\sqrt{T\gamma_{T}})$, in which multiplicative $\sqrt{\log T}$ factor is improved compared with $O(\sqrt{T\gamma_{T}\log T})$ achieved by GP-UCB (Srinivas et al., 2010) and RGP-UCB. Second, more importantly, IRGP-UCB does not need to schedule the parameters $s_{t}$ and $\lambda$ of $\zeta_{t}$ in contrast to GP-UCB and RGP-UCB. Therefore, through randomization, IRGP-UCB essentially alleviates the problem that the well-known $\beta_{t}\propto\log(t)$ strategy results in a practically too large confidence parameter. Further, over-exploration in the later iterations can be avoided. The IRGP-UCB is the first GP-UCB-based method that enjoys the above technical and practical benefits.
On the other hand, note that the dependence $\sqrt{\log|{\cal X}|}$ remains as with the prior works (Srinivas et al., 2010; Russo and Van Roy, 2014; Kandasamy et al., 2018; Paria et al., 2020). In BO, we are usually interested in the case that the total number of iterations $T<|{\cal X}|$. Thus, $\sqrt{\log|{\cal X}|}$ is dominant compared with $\sqrt{\log T}$ factor improvement.
@@ 4.2 Conditional Expected Regret Bounds
In this section, we analyze the conditional expected regret. The regret depends on the randomness of $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and $\{\zeta_{t}\}_{t\geq 1}$. Although the randomness of $f$ and $\{\epsilon_{t}\}_{t\geq 1}$ arise from the problem setup, $\{\zeta_{t}\}_{t\geq 1}$ is produced by the user. Therefore, these randomnesses should be distinguished, and we analyze the following conditional expected regret:
@@
where the expectation is taken with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$. Note that $\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}[f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}]$ is a random variable depending on $\{\zeta_{t}\}_{t\geq 1}$. Hence, showing the sub-linear high probability upper bound of the conditional expected regret implies that the BO algorithm achieves a good performance averaged over a random environment with a high probability concerning the randomness of the algorithm.
Regarding the deterministic BO algorithm, the above conditional expected regret is equivalent to the usual expected regret. Therefore, the deterministic BO algorithm with the sub-linear expected regret almost surely achieves the sub-linear conditional expected regret. On the other hand, the expected regret bound of the stochastic algorithm does not necessarily indicate the conditional expected regret bound. Thus, to compare the stochastic and the deterministic algorithms fairly, comparing not only the expected regret but also the conditional expected regret is essential.
First, we analyze the case of finite ${\cal X}$. We can obtain the following theorem:
@@
$\leq 6\sqrt{T\log(2/\delta)}+\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log(2/\delta)}+2\log(2/\delta)\right)}$
$=O\left(\sqrt{T\gamma_{T}\log|{\cal X}|+\gamma_{T}\sqrt{T\log(1/\delta)}}+\gamma_{T}\log(1/\delta)\right),$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
Since we here show the proof sketch only, See Appendix B.1 for the details. We can obtain the upper bound by the properties of the expectation and Lemma 4.2:
@@
Then, $A_{2}$ can be bounded from above using MIG and the upper bound of the chi-square random variable (Laurent and Massart, 2000). Furthermore, we can show that $A_{1}$ is a sum of the martingale sequence adapted to the history ${\cal H}_{t}=\{\zeta_{i}\}_{i\leq t}$. Therefore, we can apply Azuma’s inequality (Azuma, 1967). Finally, by combining the upper bounds of $A_{1}$ and $A_{2}$ using the union bond, we can obtain the desired result. ∎
Theorem 4.3 maintains the rate $O(\sqrt{T\gamma_{T}})$ regarding $T$. In addition, if $T>\log(1/\delta)$, the terms depending on $\delta$ can be ignored as the order. Therefore, if we compare the conditional expected regret of IRGP-UCB and the expected regret of GP-UCB, we can still observe the superiority of IRGP-UCB for the case of finite ${\cal X}$. Furthermore, the distribution of $\zeta_{t}$ is the same as in Theorem 4.1, and increasing the confidence parameter is not required.
@@
$\leq\frac{\pi^{2}}{6}+6\sqrt{T\log(2/\delta)}+\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log(2/\delta)}+2\log(2/\delta)\right)},$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
See Appendix B.2 for the proof. The resulting order maintains the same order as the expected regret bound shown in Theorem 4.2 if $T>\log(2/\delta)$.
4.3 High Probability Regret Bounds
In this section, we consider the high probability regret bounds of the randomized GP-UCB. For simplicity, we focus on the case of the finite ${\cal X}$. As discussed in (Russo and Van Roy, 2014), a direct consequence of Markov’s inequality is
@@ 4.3 High Probability Regret Bounds
Assume the same condition as in Theoremf 4.1. Then, by running IRGP-UCB, the following holds:
@@
However, compared with the existing bound of GP-UCB $O(\sqrt{T\gamma_{T}\log(T|{\cal X}|/\delta)})$, the dependence on $\delta$ is worse. Therefore, we discuss obtaining the upper bound with the dependence $\log(1/\delta)$.
If we accept increasing $\mathbb{E}[\zeta_{t}]$, we can show the following lemma:
@@
Note that this analysis can be seen as a generalized version of the analysis of GP-UCB since the randomized confidence parameter is equivalent to GP-UCB if we set $U_{t}=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ almost surely.
@@
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}$ follows a shifted exponential distribution with $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(12\delta))$ and $\lambda=1/2$ for any $t\geq 1$. Then, by running IRGP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$R_{T}$ $\leq 2\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}}+2\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}\right)}$
$=O\bigl{(}\sqrt{T\gamma_{T}\log(|{\cal X}|T/\delta)}\bigr{)}.$
See Appendix C.1 for the proof. Although Theorem 4.5 shows the sub-linear upper bounds, we can see that
$\mathbb{E}[\zeta_{t}]=\mathbb{E}[-2\log(U_{t})]\geq-2\log(\mathbb{E}[U_{t}])=2\log\left(\frac{|{\cal X}|t^{2}\pi^{2}}{6\delta}\right).$
Therefore, although $\zeta_{t}$ may be smaller than the usual $\beta_{t}$ in GP-UCB, $\mathbb{E}[\zeta_{t}]$ is larger than or equal to the usual $\beta_{t}$. We leave the high probability regret analysis without increasing the confidence parameter as future work.

### v3

@@ 3 Algorithm
Algorithm 1 shows the pseudo-code of IRGP-UCB. The difference from the original GP-UCB is that $\{\zeta_{t}\}$ is a sequence of random variables, not a monotonically increasing (non-random) sequence. Using the generated $\zeta_{t}$, the next input is chosen as follows:
@@
where $\lambda$ is a rate parameter. This can be seen as the distribution of the sum of $s_{t}$ and $Z\sim{\rm Exp}(\lambda)$. Thus, sampling from the shifted exponential distribution can be easily performed using an exponential distribution as follows:
@@ 4 Regret Analysis
In this section, we analyze the performance of IRGP-UCB and the necessity of randomizing confidence parameters. First, we analyze the average-case performance through expected regret averaged over all the randomness of $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and $\{\zeta_{t}\}_{t\geq 1}$. Second, we analyze conditional expected regret averaged over $f$ and $\{\epsilon_{t}\}_{t\geq 1}$ and conditioned on $\{\zeta_{t}\}_{t\geq 1}$. Third, we derive high-probability regret bounds of randomized GP-UCB with increasing confidence parameters. Lastly, we show that GP-UCB using a constant confidence parameter can incur linearly growing expected cumulative regret.
@@ 4.1 Expected Regret Bounds
Here, we derive the expected cumulative regret bound, which is also called Bayesian cumulative regret (BCR) (Russo and Van Roy, 2014). The definition of BCR is
@@
Suppose that $f$ is a sample path from a GP with zero mean and a positive semidefinite kernel $k$, and ${\cal X}$ is finite. Pick $\delta\in(0,1)$ and $t\geq 1$. Then, for any given ${\cal D}_{t-1}$,
@@
See Appendix A.1 for the proof. This lemma differs slightly from Lemma 5.1 of Srinivas et al. (2010), since, in Lemma 4.1, the iteration $t$ is fixed, and $\beta_{\delta}$ does not depend on $t$.
Based on Lemma 4.1, we show the following lemma, which will play a key role in the subsequent analyses:
@@
We here show a short proof of Lemma 4.2, although a detailed proof is shown in Appendix A.2. From Lemma 4.1, we can obtain the following inequality:
@@
Lemma 4.2 shows a direct upper bound of the expectation of $f(\boldsymbol{x}^{*})$. A notable consequence of Lemma 4.2 is an upper bound of BCR as follows:
@@
where $v_{t}(\boldsymbol{x})\coloneqq\mu_{t-1}(\boldsymbol{x})+\zeta^{1/2}_{t}\sigma_{t-1}(\boldsymbol{x})$. The above inequality is obtained by the selection rule of IRGP-UCB. Importantly, $\mathbb{E}[\zeta_{t}]=2+s_{t}$ is a constant since $s_{t}=2\log(|{\cal X}|/2)$ does not depend on $t$ in Lemma 4.2. Thus, using Lemma 4.2, we can obtain the BCR bound for finite input domains:
@@
Theorem 4.1 has an important implication that IRGP-UCB does not need to schedule the parameters $s_{t}$ and $\lambda$ of $\zeta_{t}$, in contrast to GP-UCB and RGP-UCB. Therefore, through randomization, IRGP-UCB essentially alleviates the problem that the well-known $\beta_{t}\propto\log(t)$ strategy results in a practically too large confidence parameter. Further, over-exploration in the later iterations can be avoided. IRGP-UCB is the first GP-UCB-based method that enjoys the above practical benefits without sacrificing regret guarantee.
On the other hand, note that the $\sqrt{\log|{\cal X}|}$ dependence remains as with the prior works (Srinivas et al., 2010; Russo and Van Roy, 2014; Kandasamy et al., 2018; Paria et al., 2020). In BO, we are usually interested in the case that the total number of iterations $T<|{\cal X}|$. Therefore, in such case, although our regret upper bound $O\bigl{(}\sqrt{\gamma_{T}T\log|{\cal X}|}\bigr{)}$ is slightly better than the existing one $O\bigl{(}\sqrt{\gamma_{T}T\log(|{\cal X}|T)}\bigr{)}$, the difference between those regret upper bounds is only the constant term.
@@ 4.2 Conditional Expected Regret Bounds
In this section, we analyze the conditional expected regret of IRGP-UCB. This analysis is newly provided compared with the previous conference version of our paper (Takeno et al., 2023). The regret depends on the randomness of $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and $\{\zeta_{t}\}_{t\geq 1}$. Although the randomness of $f$ and $\{\epsilon_{t}\}_{t\geq 1}$ arise from the problem setup, $\{\zeta_{t}\}_{t\geq 1}$ is produced by the user. Hence, distinguishing the randomness, we analyze the following conditional expected regret:
@@
where the expectation is taken with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$. Note that $\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}[f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}]$ is a random variable depending on $\{\zeta_{t}\}_{t\geq 1}$. Hence, showing the sub-linear high-probability upper bound of the conditional expected regret implies that the BO algorithm achieves a good performance, averaged over a random environment, with a high-probability concerning the randomness of the algorithm.
Let ${\cal P}$ be a set of random variables included in the problem setup (${\cal P}=\{f,\{\epsilon_{t}\}_{t\geq 1}\}$) and ${\cal A}$ be a set of random variables included in the algorithm (${\cal A}=\{\zeta_{t}\}_{t\geq 1}$ in the case of IRGP-UCB). In the usual expected regret analysis of the deterministic BO algorithm, such as GP-UCB (for example, Takeno et al., 2023; Paria et al., 2020), the expectation over the problem setup $\mathbb{E}_{{\cal P}}[R_{T}]$ is considered. When we interpret the confidence parameter in GP-UCB as a random variable equal to $\beta_{t}$ with probability 1 (${\cal A}=\{\beta_{t}\}_{t\geq 1}$), $\mathbb{E}_{{\cal P}}[R_{T}]$ is equivalent to the joint expectation $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]$. Further, we obviously see that the conditional expected regret is also equivalent: $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]=\mathbb{E}_{{\cal P}}[R_{T}\mid{\cal A}]$. On the other hand, since IRGP-UCB is stochastic, possible definitions of the expected regret are $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]$ and $\mathbb{E}_{{\cal P}}[R_{T}\mid{\cal A}]$, but these two regrets can be different. Theorems 4.1 and 4.2 show the theoretical rate of $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]$. However, these theorems do not guarantee anything about the conditional expected regret $\mathbb{E}_{{\cal P}}[R_{T}\mid{\cal A}]$. The following remark provides a motivation for deriving a conditional expected regret bound:
Remark 4.1.
Although $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]$ derived by Theorems 4.1 and 4.2 reveal average behavior of the regret of IRGP-UCB, it obscures variability caused by $\{\zeta_{t}\}_{t\geq 1}$. Hence, comparing $\mathbb{E}_{{\cal P}}[R_{T}]$ of GP-UCB and $\mathbb{E}_{{\cal P},{\cal A}}[R_{T}]$ of IRGP-UCB can be unfair. On the other hand, $\mathbb{E}_{{\cal P}}[R_{T}\mid{\cal A}]$ is a random variable, whose randomness stems from $\{\zeta_{t}\}_{t\geq 1}$. Therefore, by showing a high-probability bound of $\mathbb{E}_{{\cal P}}[R_{T}\mid{\cal A}]$, we can 1) fairly compare the regret of IRGP-UCB with $\mathbb{E}_{{\cal P}}[R_{T}]$ of GP-UCB in a sense that the expectation is taken over the same random variables ${\cal P}$, and 2) demonstrate that the expected regret of IRGP-UCB can be stably bounded with respect to the random behavior of the algorithm.
First, we analyze the case of finite ${\cal X}$. We show the following theorem:
@@
$\Pr\left(\forall T\geq 1,\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}\right]\leq U\bigl{(}T,\delta\bigr{)}\right)\geq 1-\delta,$
where
$U(T,\delta)$ $=6\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)}$
and $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
We show the proof sketch here only. See Appendix B.1 for the details.
The first key point of the proof is transforming the conditional expected regret using Lemma 4.2 and the properties of the expectation as follows:
@@
where $v_{t}(\boldsymbol{x}_{t})=\mu_{t-1}(\boldsymbol{x}_{t})+\zeta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})$. Then, $A_{2}$ can be bounded from above using MIG and the high-probability upper bound of chi-square random variables (Laurent and Massart, 2000) shown in Lemma E.4.
The second technical key point is bounding $A_{1}$ by Azuma’s inequality (Azuma, 1967), which requires the conditions that $A_{1}$ is a sum of the martingale difference sequence with conditional subgaussian property (See Lemma B.1 for details). We show that $A_{1}$ satisfies this condition using several well-known lemmas shown in Appendix B.1. Finally, we can obtain the desired result by applying the union bound for the upper bounds of $A_{1}$ and $A_{2}$ for all $T\geq 1$. ∎
Theorem 4.3 shows the same benefits as Theorem 4.1. First, under the above mild condition $\min\{s_{T}^{2}/2,s_{T}\}T>\log\bigl{(}\pi^{2}T^{2}/(3\delta)\bigr{)}$, the dominant term in $U(T,\delta)$ is $\sqrt{2C_{1}T\gamma_{T}s_{T}}$, which is the same as Theorem 4.1. Therefore, Theorem 4.3 maintains the rate $O\bigl{(}\sqrt{T\gamma_{T}\log(|{\cal X}|})\bigr{)}$ under mild condition. Second, the distribution of $\zeta_{t}$ is the same as in Theorem 4.1. Hence, increasing the confidence parameter is also avoided even if we consider the high-probability bound regarding the randomness of the algorithm.
@@
$\Pr\left(\forall T\geq 1,\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}\right]\leq U\bigl{(}T,\delta\bigr{)}\right)\geq 1-\delta,$
where
$U(T,\delta)$ $=\frac{\pi^{2}}{6}+6\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)},$
and $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
See Appendix B.2 for the proof. The resulting upper bound maintains the same rate as the expected regret bound shown in Theorem 4.2 if $\min\{s_{T}^{2}/2,s_{T}\}T>2\log\bigl{(}\pi^{2}T^{2}/(3\delta)\bigr{)}$.
4.3 High-Probability Regret Bounds
In this section, we consider the high-probability regret bounds of the randomized GP-UCB, which is newly added compared with the previous conference version of this paper. For simplicity, we focus on the case of the finite ${\cal X}$. As discussed in (Russo and Van Roy, 2014), a direct consequence of Markov’s inequality is
@@ 4.3 High-Probability Regret Bounds
Assume the same condition as in Theorem 4.1. Then, by running IRGP-UCB, for some fixed $T\geq 1$, the following holds:
@@
However, this proposition shows the regret bound for some fixed $T\geq 1$. Ideally, an anytime algorithm that guarantees that the sub-linear regret bound holds for all $T\geq 1$ with high-probability is desired. For example, the analysis of GP-UCB (Srinivas et al., 2010) is anytime and shows regret upper bound $O(\sqrt{T\gamma_{T}\log(T|{\cal X}|/\delta)})$. In addition, the resulting rate $O(1/\delta)$ regarding $\delta$ is worse than that of GP-UCB $O(\log(1/\delta))$. Therefore, we focus on obtaining an anytime regret upper bound with the $\log(1/\delta)$ dependence regarding $\delta$.
First, we consider high-probability bounds using the randomized confidence parameter. Since the randomized confidence parameter $\zeta_{t}$ is scheduled as $\mathbb{E}[\zeta_{t}]=\Theta(\log t)$, we can derive the following lemma as with Lemma 5.1 in (Srinivas et al., 2010):
@@
This analysis can be seen as a generalized version of the analysis of GP-UCB since the randomized confidence parameter is equivalent to GP-UCB if we set $U_{t}=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ almost surely.
@@
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}=2\log(1/U_{t})$, where $\{U_{t}\}_{t\geq 1}$ is a sequence of mutually independent random variables that satisfy $\mathbb{E}[U_{t}]=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ and $\Pr(U_{t}\in(0,1))=1$ for all $t\geq 1$. Then, by running the randomized GP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$\Pr\left(\forall T\geq 1,R_{T}\leq 2\sqrt{C_{1}\gamma_{T}\sum_{t=1}^{T}\zeta_{t}}\right)\geq 1-\delta,$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
See Appendix C for the proof. Note that the resulting upper bound contains the random quantity $\zeta_{t}$. Thus, this theorem still does not rigorously imply a sub-linear regret upper bound. On the other hand, we can see that Theorem 4.5 is a generalized version of the analysis of GP-UCB (Srinivas et al., 2010) since the randomized GP-UCB is equivalent to GP-UCB if $U_{t}=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ almost surely. If we adopt Theorem 4.5 to IRGP-UCB and additionally take the upper bound regarding $\sum_{t=1}^{T}\zeta_{t}$, we can obtain the following theorem:
Corollary 4.1.
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}$ follows a shifted exponential distribution with $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(6\delta))$ and $\lambda=1/2$ for any $t\geq 1$. Then, by running IRGP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$\Pr\left(\forall T\geq 1,R_{T}\leq 2\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)}\right)\geq 1-\delta,$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
See Appendix C for the proof. We can see that even if confidence parameters are randomized, the same order regret upper bound as that of GP-UCB can be obtained. Thus, this result implies that randomization does not deteriorate the high-probability regret upper bounds compared with the existing result shown by Srinivas et al. (2010).
However, in Theorem 4.5, we can see that
$\mathbb{E}[\zeta_{t}]=\mathbb{E}[-2\log(U_{t})]\geq-2\log(\mathbb{E}[U_{t}])=2\log\left(\frac{|{\cal X}|t^{2}\pi^{2}}{3\delta}\right).$
Therefore, although $\zeta_{t}$ may be smaller than the usual $\beta_{t}$ in GP-UCB, its expectation $\mathbb{E}[\zeta_{t}]$ is larger than the usual $\beta_{t}$. We leave the high-probability regret analysis without increasing the confidence parameter as future work.
4.4 Expected Regret of GP-UCB with Constant Confidence Parameter
In this section, we show an expected regret lower bound of GP-UCB with any constant confidence parameter. This analysis is added from the previous conference version of this paper. We can show that GP-UCB, using any constant confidence parameter, incurs linear expected cumulative regret in the following simple problem instance:
Theorem 4.6.
Assume that ${\cal X}=\{\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}\}$, $\epsilon_{t}\sim{\cal N}(0,1)$ for all $t\geq 1$, and
$\left(\begin{array}[]{c}f(\boldsymbol{x}^{(1)})\\ f(\boldsymbol{x}^{(2)})\end{array}\right)\sim{\cal N}\left(\left(\begin{array}[]{c}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}1&\rho\\ \rho&0.99\end{array}\right)\right),$
where $\rho<1$ is a covariance parameter. Then, if GP-UCB with any constant confidence parameter $\beta_{t}^{(1/2)}=c>0$ runs, then BCR grows linearly, that is,
${\rm BCR}_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\right]=\Omega(T).$
See Appendix D for the proof. In the proof, we show that the probability of the extreme case, where the GP-UCB algorithm continue to evaluate $\boldsymbol{x}^{(1)}$ even if $f(\boldsymbol{x}^{(1)})<f(\boldsymbol{x}^{(2)})$, is rigorously positive.
Theorem 4.6 suggests that, under assumptions of Theorems 4.1 and 4.3, the confidence parameter must be increased in the GP-UCB algorithm to obtain sub-linear expected regret upper bounds. Therefore, randomizing confidence parameters as in IRGP-UCB is necessary to avoid an increase in confidence parameters.